when you add 18 to 1/4 of a number you get the number itself
Answer by rfer(16322)

1/4x+18=x
1/4x-x=-18
-3/4x=-18
x=-18/(-3/4)
x=-18*-4/3
x=24
--------------
1/4*24+18=24